OpenStudy (anonymous):

What is the 6th term of the geometric sequence where a1 = 625 and a2 = -125?

OpenStudy (anonymous):

these is a result of using \[a_n=a_1r^{n-1}\]

OpenStudy (anonymous):

\[a_6=a_1r^5=625r^5\] so for r we use\[a_2=a_1r^{2-1}=625r=-125\] so now solve for r then place back in a_6 equation

OpenStudy (anonymous):

r=-125/625=-1/5 \[a_6=625r^5=625(-1/5)^5\]